OpenStudy (anonymous):

2x + 1 greater than or equal to 5 or -3x > -9

OpenStudy (anonymous):

@jdoe0001

OpenStudy (gorv):

2x+1>=5 -3x>-9 2x>=4 3x<9 x>=4/2 x<3 x>=2

OpenStudy (gorv):

\[\in[2,3)\]

OpenStudy (gorv):

@BlveXO

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Can you also explain this one please?

OpenStudy (anonymous):

[5.01] -2y - 7 > 11

OpenStudy (anonymous):

@gorv